OpenStudy (anonymous):

use multiplying by 1 to find an expression equivelent to 11/18 with a denominator of 18z. The equivelent expression is?

OpenStudy (radar):

z/z=1 (1/18) times 1 is still 1/18 does not change the value now (1/18) times z/z = z/18z which still has the value of 1/18

OpenStudy (radar):

I know too easy, but an important concept.

OpenStudy (radar):

do you understand?

OpenStudy (anonymous):

yes, i kinda do understand.

OpenStudy (radar):

do you understand that anything divided by it self is 1 ?

OpenStudy (anonymous):

yes

OpenStudy (radar):

then z/z or z divided by itself is 1. Right?

OpenStudy (anonymous):

yes its 1

OpenStudy (radar):

Now for the last piece. Do you undertstand when you multiply anything by 1 you do not change its value?

OpenStudy (anonymous):

yes i understand that too

OpenStudy (anonymous):

its going to be 11/18

OpenStudy (anonymous):

becauese i am mulltiplying by 1

OpenStudy (radar):

we multiplied 1/18 by 1 because z/z is 1 ok we multiplied 11/18 by 1 or z/z so it will now look like 11z/18z I thought we were working with 1/18. Still we haven't changed it its value, just its appearance.

OpenStudy (radar):

Sorry, I slipped up, no wonder you kinda understood it lol

OpenStudy (radar):

better now huh?

OpenStudy (anonymous):

yes :)

OpenStudy (radar):

\[11z \over18z\]

OpenStudy (radar):

Good luck.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

11 -- 18